1. An unstable particle with mass 2.81 x 10-27 kg is initially at rest. The particle decays into two fragments that fly off awhere the relativistic gamma factor is defined by r- y is always a number greater than one.
Answer #1

At a time,4 parts can be solved.Still these will help you in others too.Comment regarding any query .Thanks...
Answer #2

(a) There is not any force is acting on the particles so the system is isolated.

(b) For the analysis of this situation we can use 1. law of conservation of momentum and 2. law of conservation of energy.

(c)  

  \gamma_1=\frac{1}{\sqrt{1-\frac{u_1^2}{c^2}}}

here u_1=0.987 c , put this value in above equation we will get

\gamma_1=6.22

similarly

\gamma_2=\frac{1}{\sqrt{1-\frac{u_2^2}{c^2}}}

u_2=-0.898c

we get

72 = 2.27   

(d)